# Minimax optimal differentially private synthetic data for smooth queries

### v1

@@ 4.2 Accuracy
    $\mathbb{E}d_{k}(\mu_{{X}},q)$ $\leq\sqrt{d}\cdot\Delta+\frac{2C_{k}^{\mathrm{Jac}}\cdot d}{m^{k}}+\sqrt{C_{k}}\cdot\frac{c^{d}\sqrt{\log(1.25/\delta)}}{\varepsilon n}\cdot 2$
@@
Here we can take $C_{k}^{\prime}=2C_{k}^{\mathrm{Jac}}c^{k}+{2\sqrt{C_{k}}\cdot c^{k}}+1\leq C^{k}{k!}$ for some absolute constant $C>0$.
@@ Appendix A Minimax lower bound
A.1 Total variance distance bound
@@ A.2 Construction of a family of test functions
$\chi({\mathbf{u}}):=u_{1}^{k}\prod_{\ell=1}^{d}\eta(u_{\ell}),\qquad{\mathbf{u}}=(u_{1},\dots,u_{d})\in\mathbb{R}^{d},$
@@
$C_{0}:=\max_{1\leq|\alpha|\leq k}\big\|\partial^{\alpha}\chi\big\|_{\infty}\in(0,\infty).$
From our construction of $\chi({\mathbf{u}})$, $C_{0}$ is a constant depending only on $k$. Then, for each $t\in[M]$, define
@@
(ii) bounds on derivatives For any multi-index $\alpha$ with $1\leq|\alpha|\leq k$, by the chain rule,
$\partial^{\alpha}f_{t}({\mathbf{x}})=\frac{r^{k-|\alpha|}}{C_{0}}\,(\partial^{\alpha}\chi)\!\left(\frac{{\mathbf{x}}-{\mathbf{a}}_{t}}{r}\right).$
@@
$\|\partial^{\alpha}f_{t}\|_{\infty}\leq\frac{r^{k-|\alpha|}}{C_{0}}\,\|\partial^{\alpha}\chi\|_{\infty}\leq\frac{1}{C_{0}}\max_{|\beta|\leq k}\|\partial^{\beta}\chi\|_{\infty}=1,$
where we used $r\leq 1$ and $k-|\alpha|\geq 0$. Therefore $f_{t}\in\mathcal{F}_{k}$.
@@
$f_{t}({\mathbf{x}}_{t}^{\max})=\frac{r^{k}}{C_{0}}\chi({\mathbf{u}}^{\star})=\frac{r^{k}}{C_{0}}\Big(\frac{1}{16}\Big)^{k}.$
Finally, setting $c^{\max}_{k}=(16^{-k})/C_{0}>0$ gives $f_{t}({\mathbf{x}}_{t}^{\max})\geq c^{\max}_{k}r^{k}$. ∎
@@ A.3 Construction of input datasets
and $v_{i}=\operatorname{TV}\big(Q_{{(i,+)}},Q_{{(i,+)}}\big)$. Let $(Z_{(i,+)},Z_{(i,-)})$ be a maximal coupling of $Q_{{(i,+)}}$ and $Q_{{(i,-)}}$ so that
@@ A.4 Proof of Theorem 2
Proof of Theorem 2.
Let $\mathcal{M}$ be any $(\varepsilon,\delta)$–DP algorithm that outputs synthetic data in $1,1]^{d}$. Let $\theta\sim\mathrm{Unif}(\{\pm 1\}^{M})$, $M=m^{d}$, $Z\coloneqq\mathcal{M}(X^{\theta})$, and $p_{Z}$ be its associated empirical distribution. Let $Q_{(i,+)}$ and $Q_{(i,-)}$ be the distribution of $Z$ conditioning on the events $\{\theta_{i}=+1\}$ and $\{\theta_{i}=-1\}$, correspondingly, and let
@@ Appendix C Proof of Theorem 3
To show the infinity norm bound in Lemma 9, we introduce the Bernstein inequality in harmonic analysis, which states that the infinity norm of a triangular polynomial can be controlled by its derivative.
Lemma 10 (A Bernstein inequality for trigonometric polynomials, [43], Ex. 7.15(c), (7.25)).
Let $Q(\vartheta)=\sum_{n=-m}^{m}a_{n}e^{in\vartheta}$ be a trigonometric polynomial of degree $\leq m$. Then
@@
Proof of Lemma 9.
Induct on $|{\boldsymbol{\alpha}}|$. For ${\boldsymbol{\alpha}}=0$, we have $P_{0,0}\equiv 1$.
@@
Proof of Theorem 3.
We will show
@@ Appendix D Multivariate Jackson’s Theorem for $k$-smooth functions
    (Triangular polynomial) $J_{1}(x)=\sum_{i=0}^{m}\widehat{J}_{1}(i)\cos(ix)$;
@@
Proof of Lemma 3.
For any even periodic function $G({\boldsymbol{\theta}})=G(\theta_{1},\dots,\theta_{d})$, define operator $\mathcal{A}_{i}$ such that $\mathcal{A}_{i}(G)$ is to apply Lemma 12 on the variable $\theta_{i}$:
@@
Taking maximum over $\boldsymbol{\alpha}$} concludes the proof. ∎
Appendix E Estimation for $S$.
@@
$\mathcal{A}_{j}:=\bigl\{K\in\{0,\ldots,m\}^{d}:\ \|K\|_{\infty}=j\bigr\},\qquad\|K\|_{\infty}:=\max_{1\leq i\leq d}|K_{i}|.$
@@
Appendix F Computational complexity of Algorithm 1

### v2

@@ 4.2 Accuracy
    $\mathbb{E}d_{k}(p_{{X}},p_{Y})$ $\leq d\Delta+\frac{2C_{k}^{\mathrm{Jac}}\cdot d}{m^{k}}+\sqrt{C_{k}}\cdot\frac{c^{d}\sqrt{\log(1.25/\delta)}}{\varepsilon n}\cdot 2+k(\varepsilon n)^{-1}$
@@
Here we can take $C_{k}^{\prime}=2C_{k}^{\mathrm{Jac}}c^{k}+{2\sqrt{C_{k}}\cdot c^{k}}+1\leq C^{k}{k!}$ for some absolute constant $C>0$. Moreover, from (7), with the choice of $m$ above, we have
[…]
This finishes the proof.
@@ Appendix A Minimax lower bound
A.1 Total variation distance bound
@@ A.2 Construction of a family of test functions
$\chi({\mathbf{u}}):=u_{1}\prod_{\ell=1}^{d}\eta(u_{\ell}),\qquad{\mathbf{u}}=(u_{1},\dots,u_{d})\in\mathbb{R}^{d},$
@@
$C_{0}:=\max_{1\leq|{\boldsymbol{\alpha}}|\leq k}\big\|\partial^{\boldsymbol{\alpha}}\chi\big\|_{\infty}\in(0,\infty).$
From our construction of $\chi({\mathbf{u}})$, $C_{0}$ is a constant depending only on $k$. (It suffices to consider the case $k<d$. Then we are taking derivatives to at most $k$ many $\eta(u_{\ell})$, and we can upper bound the rest by $|\eta(u_{\ell^{\prime}})|\leq 1$ for $\ell^{\prime}\neq\ell$.)
For each $t\in[M]$, define
@@
(ii) bounds on derivatives For any multi-index $\boldsymbol{\alpha}$} with $1\leq|{\boldsymbol{\alpha}}|\leq k$, by the chain rule,
$\partial^{\boldsymbol{\alpha}}f_{t}({\mathbf{x}})=\frac{r^{k-|{\boldsymbol{\alpha}}|}}{C_{0}}\,(\partial^{\boldsymbol{\alpha}}\chi)\!\left(\frac{{\mathbf{x}}-{\mathbf{a}}_{t}}{r}\right).$
@@
$\|\partial^{\boldsymbol{\alpha}}f_{t}\|_{\infty}\leq\frac{r^{k-|{\boldsymbol{\alpha}}|}}{C_{0}}\,\|\partial^{\boldsymbol{\alpha}}\chi\|_{\infty}\leq\frac{1}{C_{0}}\max_{1\leq|{\boldsymbol{\beta}}|\leq k}\|\partial^{\boldsymbol{\beta}}\chi\|_{\infty}=1,$
where we used $r\leq 1$ and $k-|{\boldsymbol{\alpha}}|\geq 0$. Therefore $f_{t}\in\mathcal{F}_{k}$.
@@
$f_{t}({\mathbf{x}}_{t}^{\max})=\frac{r^{k}}{C_{0}}\chi({\mathbf{u}}^{\star})=\frac{r^{k}}{16C_{0}}.$
Finally, setting $c^{\max}_{k}=1/(16C_{0})>0$ gives $f_{t}({\mathbf{x}}_{t}^{\max})\geq c^{\max}_{k}r^{k}$. ∎
@@ A.3 Construction of input datasets
and $v_{i}=\operatorname{TV}\big(Q_{{(i,+)}},Q_{{(i,-)}}\big)$. Let $(Z_{(i,+)},Z_{(i,-)})$ be a maximal coupling of $Q_{{(i,+)}}$ and $Q_{{(i,-)}}$ so that
@@ A.4 Proof of Theorem 2
Proof.
(Proof of Theorem 2) Let $\mathcal{M}$ be any $(\varepsilon,\delta)$–DP algorithm that outputs synthetic data in $1,1]^{d}$. Let $\theta\sim\mathrm{Unif}(\{\pm 1\}^{M})$, $M=m^{d}$, $Z\coloneqq\mathcal{M}(X^{\theta})$, and $p_{Z}$ be its associated empirical distribution. Let $Q_{(i,+)}$ and $Q_{(i,-)}$ be the distribution of $Z$ conditioning on the events $\{\theta_{i}=+1\}$ and $\{\theta_{i}=-1\}$, correspondingly, and let
@@ Appendix C Proof of Theorem 3
To show the infinity norm bound in Lemma 9, we introduce the Bernstein inequality in harmonic analysis, which states that the infinity norm of a trigonometric polynomial can be controlled by its derivative.
Lemma 10 (Bernstein inequality [43] Ex. 7.15(c), (7.25)).
Consider a trigonometric polynomial $Q(\vartheta)=\sum_{n=-m}^{m}a_{n}e^{in\vartheta}$ of degree $\leq m$. Then
@@
Proof.
(Proof of Lemma 9) Induct on $|{\boldsymbol{\alpha}}|$. For ${\boldsymbol{\alpha}}=0$, we have $P_{0,0}\equiv 1$.
@@
Proof.
(Proof of Theorem 3) We will show
@@ Appendix D Multivariate Jackson’s Theorem for $k$-smooth functions
    (Trigonometric polynomial) $J_{1}(x)=\sum_{i=0}^{m}\widehat{J}_{1}(i)\cos(ix)$;
@@
Proof.
(Proof of Lemma 3) For any even periodic function $G({\boldsymbol{\theta}})=G(\theta_{1},\dots,\theta_{d})$, define operator $\mathcal{A}_{i}$ such that $\mathcal{A}_{i}(G)$ is to apply Lemma 12 on the variable $\theta_{i}$:
@@
Taking the maximum over $\boldsymbol{\alpha}$} concludes the proof. ∎
Appendix E Proof of Theorem 4
Proof.
For any function $f\in\mathcal{F}_{k}$, let $\tilde{f}=\sum_{K}\tilde{c}_{K}\overline{T}_{K}$ be the multivariate Chebyshev polynomial in in Lemma 3. We have
Appendix F Estimation for $S$.
@@ Appendix E Proof of Theorem 4
$\mathcal{A}_{j}:=\bigl\{K\in\{0,\ldots,m\}^{d}\setminus\{\mathbf{0}\}:\ \|K\|_{\infty}=j\bigr\},\qquad\|K\|_{\infty}:=\max_{1\leq i\leq d}|K_{i}|.$
@@
Appendix G Computational complexity of Algorithm 1